# Smoothed Graph Contrastive Learning via Seamless Proximity Integration

### v1

@@ Abstract
Graph contrastive learning (GCL) aligns node representations by classifying node pairs into positives and negatives using a selection process that typically relies on establishing correspondences within two augmented graphs. The conventional GCL approaches incorporate negative samples uniformly in the contrastive loss, resulting in the equal treatment negative nodes, regardless of their proximity to the true positive. In this paper, we present a Smoothed Graph Contrastive Learning model (SGCL), which leverages the geometric structure of augmented graphs to inject proximity information associated with positive/negative pairs in the contrastive loss, thus significantly regularizing the learning process. The proposed SGCL adjusts the penalties associated with node pairs in the contrastive loss by incorporating three distinct smoothing techniques that result in proximity aware positives and negatives. To enhance scalability for large-scale graphs, the proposed framework incorporates a graph batch-generating strategy that partitions the given graphs into multiple subgraphs, facilitating efficient training in separate batches. Through extensive experimentation in the unsupervised setting on various benchmarks, particularly those of large scale, we demonstrate the superiority of our proposed framework against recent baselines.
Machine Learning, ICML
@@ 1 Introduction
Graph Neural Networks (GNNs) (Gilmer et al., 2017; Kipf & Welling, 2017; Xu et al., 2019b) have developed rapidly by providing the powerful frameworks for the analysis of graph-structured data. A significant portion of GNNs primarily focus on (semi-)supervised learning, which requires access to abundant labeled data (Veličković et al., 2018; Kipf & Welling, 2017; Behmanesh et al., 2023). However, labeling graphs is challenging because they often represent specialized concepts within domains like biology.
Graph Contrastive Learning (GCL), as a new paradigm of Self-Supervised Learning (SSL) (Liu et al., 2023) in the graph domain, has emerged to address the challenge of learning meaningful representations from graph-structured data (Wu et al., 2023; Xie et al., 2023). They leverage the principles of self-supervised learning and contrastive loss (Li et al., 2019) to form a simplified representation of graph-structured data without relying on supervised data.
In a typical GCL approach, several graph views are generated through stochastic augmentations of the input graph. Subsequently, representations are learned by comparing congruent representations of each node, as an anchor instance, with its positive/negative samples from other views (Veličković et al., 2019; Zhu et al., 2020; Hassani & Khasahmadi, 2020). More specifically, the GCL approach initially captures the inherent semantics of the graph to identify the positive and negative nodes. Then, the contrastive loss efficiently pulls the representation of the positive nodes or subgraphs closer together in the embedding space while simultaneously pushing negative ones apart.
Conventional GCL methods follow a straightforward principle when distinguishing between positive and negative pairs: pairs of corresponding points in augmented views are considered positive pairs (similar), while all other pairs are regarded as negative pairs (dissimilar) (Zhu et al., 2020). This strategy ensures that for each anchor node in one augmented view, there exists one positive pair, while all remaining nodes in the second augmented view are paired as negatives.
@@
We evaluate the SGCL framework in node and graph classification tasks across benchmarks of various properties. Our results consistently demonstrate the superior performance of our method compared to state-of-the-art GCL methods. Our contributions are summarized as follows:
  • We propose a novel graph contrastive loss function that seamlessly integrates node proximity information, overcoming the uniform negative sampling limitations found in conventional GCL methods
  • We introduce three formulations for integrating proximity information into the contrastive learning loss to generate positive and negative pairs.
@@
  • We evaluate the SGCL model for both node and graph classification on various benchmarks of different scales, demonstrating its superior performance over state-of-the-art methods.
@@ 2.1 Uniform negative sampling
In the unsupervised GCL models introduced with two views $i$ and $j$, for each anchor node $v_{t}^{(i)}$ with feature embedding $\mathbf{h}_{t}^{(i)}$, the contrasting learning model defines a positive set $\mathcal{P}(v_{t}^{(i)})=\{v_{p}^{(j)}\}_{p=1}^{P}$, consisting of $P$ elements, and a negative set $\mathcal{Q}(v_{t}^{(i)})=\{v_{q}^{(j)}\}_{q=1}^{Q}$, comprising $Q$ elements.
In the absence of labeled information, these sets are confined to containing consistent samples within each graph view. Essentially, the positive set is formed by pairing embeddings in the two augmented graph views that align with the same node. Therefore, $P=1$ and $\mathcal{P}(v_{t}^{(i)})=\{v_{p}^{(j)}\}$, where $v_{p}^{(j)}$ in view $j$ corresponds to $v_{t}^{(i)}$ in view $i$. If corresponding nodes are in the same order in two views, then $t=p$. Additionally, all incongruent samples in each view $j$ are categorized as negative samples, $\mathcal{Q}(v_{t}^{(i)})=\{v_{q}^{(j)}\}_{q=1,q\neq t}^{N-1}$.
Considering the ground truth, positive/negative pairs demonstrate semantic congruence/incongruence, particularly in relation to shared labels with the anchor. These pairs encompass samples affiliated with either the same class (positive) or different classes (negative). However, in the absence of labeled information, numerous incongruent nodes are false negative because they have the potential to be semantically similar to the anchor node, as demonstrated in (Xia et al., 2021), but are instead categorized as negative pairs.
This misalignment of the negative pairs adversely affects the learning process due to its inadvertent impact on the loss function. Consider the contrastive loss function 1 for each anchor node $v_{t}^{(i)}$. The objective is to minimize the distance between embeddings of positive pair $\{v_{t}^{(i)},v_{t}^{(i)}\}$ and simultaneously maximize the distance between embeddings of negative pairs $\{v_{t}^{(i)},v_{q}^{(j)}\}_{q=1,q\neq t}^{N-1}$:
$\mathcal{L}_{con}(v_{t}^{(i)},V^{(j)})=$ (1)
$-\log{\left(\frac{\exp{(\mathbf{h}_{t}^{(i)}.\mathbf{h}_{t}^{(j)}/\tau)}}{\exp{(\mathbf{h}_{t}^{(i)}.\mathbf{h}_{t}^{(j)}/\tau)}+\sum_{q=1,q\neq t}^{N-1}\exp{(\mathbf{h}_{t}^{(i)}.\mathbf{h}_{q}^{(j)}/\tau)}}\right)}$
@@ 2.2 Motivation and intuition
Our high-level intuition involves a smoothed contrastive learning approach that leverages the inherent geometric information within a graph to assign lower penalties for the negatives that are in close proximity to the ground truth positive. As such it promotes predictions that are (similarly to conventional CL) either exactly at the ground truth positive, or (differently from conventional CL) at least in the geodesic vicinity of the positive. By introducing this information, we strongly regularize the learning process, thereby improving the overall accuracy.
In the following, we will demonstrate how leveraging the inherent geometric information within a graph can provide additional insights and enhance the performance of the GCL models.
@@ 2.3 Leveraging the advantages of graph geometry
In conventional contrastive learning models, the positive pairs between two views are represented by a positive matrix $\Pi_{pos}^{(i,j)}\in{\{0,1\}}^{N\times N}$ with ’1’ on the diagonal and a negative matrix $\Pi_{neg}^{(i,j)}=1-\Pi_{pos}^{(i,j)}\in{\{0,1\}}^{N\times N}$ with ’0’ on the diagonal and ’1’ in the off-diagonal elements.
@@
We propose a smoothing strategy that goes beyond simple binary categorization of matrices as positive or negative and applies a form of smoothing to the standard contrastive loss. This strategy allows nodes initially categorized as positive or negative to have values ranging from ’0’ to ’1’, indicating their degree of association with positive or negative samples, respectively. The smoothing process $\mathcal{S}({\Pi}_{pos}^{(i,j)},\mathbf{A}^{(i)})$ effectively integrates the geometric structure of a graph $\mathcal{G}^{(i)}$, enriching both positive and negative pairs by encompassing neighborhood relationships and capturing the broader context of the nodes $\mathcal{V}^{(i)}$.
@@
Smoothing involves iteratively updating node values based on the values of their neighboring nodes. This process helps to smooth the data while preserving the underlying graph structure. In a general setting, we take a binary positive matrix $\Pi_{pos}^{(i,j)}$ as the input of the smoothing approach $\mathcal{S}({\Pi}_{pos}^{(i,j)},\mathbf{A}^{(i)})$ to generate a smooth positive matrix $\tilde{\Pi}_{pos}^{(i,j)}\in{[0,1]}^{N\times N}$. The corresponding smoothed negative matrix $\tilde{\Pi}_{neg}^{(i,j)}$ is then obtained as $1-\tilde{\Pi}_{pos}^{(i,j)}$. In the following, we formulate three existing formulations for prompting smoothing, including Taubin smoothing (Taubin, 2023), Bilateral smoothing Tomasi & Manduchi (1998), and Diffusion-based smoothing (Gerig et al., 1992).
@@
Taubin smoothing $\mathcal{S}_{T}(\mathbf{V},\mathbf{L},K,\mu,\tau)$ involves iteratively performing two stages of filtering utilized graph Laplacian matrix $\mathbf{L}\in\mathbb{R}^{N\times N}$ to smooth the binary matrix $\mathbf{V}\in\{0,1\}^{N\times D}$ as follows:
$\mathbf{V}^{(k+1)}=(\mathbf{I}+\mu\mathbf{L})((\mathbf{I}+\tau\mathbf{L})\mathbf{V}^{(k)})$ (2)
The interior filter refers to the positive Laplacian filter, which operates by smoothing the values within the matrix $\mathbf{V}$ by moving each vertex to the average position of its neighbors. This process tends to blur the details, where $\tau>0$ is a positive constant that controls the amount of smoothing. Subsequently, the negative Laplacian filter is employed to rectify the oversmoothing that occurred in the previous step. It moves each vertex in the opposite direction, preserving important geometric features. Here, $\mu(<-\tau)$ is a negative constant that corrects the oversmoothing from the previous step.
@@
Bilateral smoothing $\mathcal{S}_{B}(\mathbf{V},\mathbf{A},\sigma_{spa},\sigma_{int})$ performs smoothing on a binary matrix $\mathbf{V}$ by combining information from nearby nodes, considering both spatial proximity and intensity similarity. Spatial proximity $d_{spa}$ captures the structural proximity between nodes within the graph by employing shortest path distance metrics. Intensity similarity $d_{int}$ is determined by evaluating the similarity in binary values between two nodes, typically quantified using metrics like the Hamming distance. The bilateral filter weight $w(i,j)$ is then computed, incorporating both criteria:
$w(i,j)=\exp{\left(-\frac{d_{spa}}{2\sigma_{spa}^{2}}-\frac{d_{int}}{2\sigma_{int}^{2}}\right)},$ (3)
where $\sigma_{spa}^{2}$ and $\sigma_{int}^{2}$ modulate the extent of both spatial and intensity smoothing, respectively. The smoothed value of the central node $v_{i}$ is calculated as a weighted average of the binary values in row $i$ based on the calculated weights:
$\tilde{\textbf{v}}_{i}=\frac{\sum_{k=1}^{N}w(i,k)\textbf{v}_{k}}{\sum_{k=1}^{N}w(i,k)}$ (4)
@@
Diffusion-based smoothing $\mathcal{S}_{D}(\mathbf{V},\mathbf{A},K,\eta)$ represents a method that employs diffusion equation to effectively diffuse information among nodes within a graph. It describes how data or attributes propagate from one node to its neighboring nodes over time, thereby achieving binary value smoothing. The smoothing process is first initialized with the original map matrix as the initial conditions, where each binary value serves as the starting ”heat” at its respective node. Subsequently, the new value for each node is iteratively calculated based on the chosen diffusion equation and the binary values of its neighbors, considering the graph’s connectivity and the diffusion process as $\textbf{v}_{i}^{(k+1)}=\textbf{v}_{i}^{(k)}+\eta\bar{\textbf{v}}_{i}^{(k)}$, where $\bar{\textbf{v}}_{i}^{(k)}=\sum_{j\in\mathcal{N}(v_{i})}\textbf{v}_{j}^{(k)}$ and a diffusion rate $\eta$ is applied to determine how much the binary value diffuses from one node to another.
It is worth noting that in the last step of all smoothing approaches, a masking step is applied to preserve the initial ’1’ values while ensuring that the other smoothed values remain unchanged.
@@
Figure 1 illustrates an example of the efficacy of the smoothing approaches. We take a specific graph, such as a grid graph, and randomly establish a delta function, centered on specific vertices within this graph, resulting in the creation of a binary matrix. Subsequently, we employ a variety of smoothing techniques on this binary matrix. Given the uniform neighborhood structure of the grid, the resulting output exhibits a Gaussian-like distribution, which its center aligned to the initial vertex. However, the varied values in the smoothed matrix are indicative of the distinct strategies employed in the smoothing process.
@@
Figure 2: In the general context of conventional contrastive learning approaches, for every anchor node $v_{4}^{(i)}$ in $\mathcal{G}^{(i)}$, a corresponding positive node $v_{4}^{(j)}$ exists in $\mathcal{G}^{(j)}$, with all other node pairs being negative (left image). Smoothing techniques, which leverages the geometry of graph $\mathcal{G}^{(i)}$, effectively extract neighboring node information of an anchor node $v_{4}^{(i)}$ and generate smoothed positive and negative pairs matrices $\tilde{\Pi}_{pos}^{(i,j)}$ and $\tilde{\Pi}_{neg}^{(i,j)}$ (right image).

### v2

@@ Abstract
Graph contrastive learning (GCL) aligns node representations by classifying node pairs into positives and negatives using a selection process that typically relies on establishing correspondences within two augmented graphs. The conventional GCL approaches incorporate negative samples uniformly in the contrastive loss, resulting in the equal treatment of negative nodes, regardless of their proximity to the true positive. In this paper, we present a Smoothed Graph Contrastive Learning model (SGCL), which leverages the geometric structure of augmented graphs to inject proximity information associated with positive/negative pairs in the contrastive loss, thus significantly regularizing the learning process. The proposed SGCL adjusts the penalties associated with node pairs in contrastive loss by incorporating three distinct smoothing techniques that result in proximity-aware positives and negatives. To enhance scalability for large-scale graphs, the proposed framework incorporates a graph batch-generating strategy that partitions the given graphs into multiple subgraphs, facilitating efficient training in separate batches. Through extensive experimentation in the unsupervised setting on various benchmarks, particularly those of large scale, we demonstrate the superiority of our proposed framework against recent baselines. The implementation is available at https://github.com/maysambehmanesh/SGCL.
@@ 1 Introduction
Graph Neural Networks (GNNs) [1, 2, 3] have developed rapidly by providing powerful frameworks for the analysis of graph-structured data. A significant portion of GNNs primarily focus on (semi-) supervised learning, which requires access to abundant labeled data [4, 2, 5]. However, labeling graphs is challenging because they often represent specialized concepts within domains like biology.
Graph Contrastive Learning (GCL), as a new paradigm of Self-Supervised Learning (SSL) [6] in the graph domain, has emerged to address the challenge of learning meaningful representations from graph-structured data [7, 8]. They leverage the principles of self-supervised learning and contrastive loss [9] to form a simplified representation of graph-structured data without relying on supervised data.
In a typical GCL approach, several graph views are generated through stochastic augmentations of the input graph. Subsequently, representations are learned by comparing congruent representations of each node, as an anchor instance, with its positive/negative samples from other views [10, 11, 12]. More specifically, the GCL approach initially captures the inherent semantics of the graph to identify the positive and negative nodes. Then, the contrastive loss efficiently pulls the representation of the positive nodes or subgraphs closer together in the embedding space while simultaneously pushing negative ones apart.
Conventional GCL methods follow a straightforward principle when distinguishing between positive and negative pairs: pairs of corresponding points in augmented views are considered positive pairs (similar), while all other pairs are regarded as negative pairs (dissimilar) [11]. This strategy ensures that for each anchor node in one augmented view, there exists one positive pair, while all remaining nodes in the second augmented view are paired as negatives.
@@
Our contributions are summarized as follows:
  • We introduce three formulations for integrating proximity information into the contrastive learning loss, aimed at improving the assignment of positive and negative pairs.
  • We devise three novel schemes for a graph contrastive loss function (i.e., SGCL-T, SGCL-B, and SGCL-D) that seamlessly integrates node proximity information, overcoming the uniform negative sampling limitations found in conventional GCL methods.
@@
  • We perform an analytical study, complemented by extensive empirical evaluations for both node and graph classification on various benchmarks, demonstrating the consistent improvement of SGCL over state-of-the-art GCL methods.
@@ 2 Background and motivation
2.1 Preliminaries
In the domain of unsupervised graph representation learning, we introduce an undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ constitutes the node set $\{{v}_{1},{v}_{2},...,{v}_{N}\}$, and $\mathcal{E}$ denotes the edge set, formally captured as $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$. Within this contextual framework, we establish the definition of two pivotal matrices: the feature matrix $\mathbf{X}\in\mathbb{R}^{N\times F}$, wherein each $\mathbf{x}_{i}\in\mathbb{R}^{F}$ represents the feature vector associated with a distinct node ${v}_{i}$; and the binary adjacency matrix $\mathbf{A}\in\{0,1\}^{N\times N}$.
Definition 2.1 (Positive and negative set).
In the context of the conventional GCL approach with two graph views $\mathcal{G}^{(i)}$ and $\mathcal{G}^{(j)}$, considering an anchor node $v_{t}^{(i)}$ in view $i$, the positive set consists of embeddings $v_{p}^{(j)}$ in view $j$ that correspond to the same node as $v_{t}^{(i)}$. Formally, this is expressed as $\mathcal{P}(v_{t}^{(i)})=\{v_{p}^{(j)}\}_{p=1}^{P}$, where $P=1$ because there is only one $v_{p}^{(j)}$ that corresponds to the $v_{t}^{(i)}$. Similarly, the negative set for $v_{t}^{(i)}$ includes all embeddings $v_{q}^{(j)}$ in view $j$ that do not correspond to the same node as $v_{t}^{(i)}$, formally expressed as $\mathcal{Q}(v_{t}^{(i)})=\{v_{q}^{(j)}\}_{q=1,q\neq t}^{Q}$, where $Q=N-1$ and $N$ denotes the total number of samples in view $j$.
@@ 2.2 Uniform negative sampling
Considering the ground truth, positive/negative pairs demonstrate semantic congruence/incongruence, particularly in relation to shared labels with the anchor. These pairs encompass samples affiliated with either the same class (positive) or different classes (negative). Nevertheless, in the absence of labeled information, numerous incongruent nodes are inevitably categorized as false negatives, even when they may share semantic similarities with the anchor node [13].
This misalignment of the negative pairs adversely affects the learning process due to its inadvertent impact on the loss function. Consider the InfoNCE contrastive loss function [14] for each anchor node $v_{t}^{(i)}$. The objective is to minimize the distance between embeddings of positive pair $\{v_{t}^{(i)},v_{t}^{(j)}\}$ and simultaneously maximize the distance between embeddings of negative pairs $\{v_{t}^{(i)},v_{q}^{(j)}\}_{q=1,q\neq t}^{N-1}$:
$\mathcal{L}_{\text{InfoNCE}}(v_{t}^{(i)},V^{(j)})=-\log{\left(\frac{\exp{(\mathbf{h}_{t}^{(i)}.\mathbf{h}_{t}^{(j)}/\tau)}}{\exp{(\mathbf{h}_{t}^{(i)}.\mathbf{h}_{t}^{(j)}/\tau)}+\sum\nolimits_{q=1,q\neq t}^{N-1}\exp{(\mathbf{h}_{t}^{(i)}.\mathbf{h}_{q}^{(j)}/\tau)}}\right)}$ (1)
@@ 2.3 Motivation and intuition
Our high-level intuition involves a smoothed contrastive learning approach that leverages the inherent geometric information within a graph to assign lower penalties for the negatives that are in close proximity to the ground truth positive. As such it promotes predictions that are (similarly to conventional CL) either exactly at the ground truth positive, or (differently from conventional CL) at least in the geodesic vicinity of the positive. By introducing this information, we strongly regularize the learning process, thereby improving the overall accuracy. In the following, we will demonstrate how leveraging the inherent geometric information within a graph can provide additional insights and enhance the performance of the GCL models.
@@ 2.4 Leveraging the advantages of graph geometry
In conventional contrastive learning models, the positive pairs between two views are represented by a positive matrix $\Pi_{\text{pos}}^{(i,j)}\in{\{0,1\}}^{N\times N}$, where the diagonal elements are ’1’ and the off-diagonal elements are ’0’. The corresponding negative matrix is defined as $\Pi_{\text{neg}}^{(i,j)}=1-\Pi_{\text{pos}}^{(i,j)}\in{\{0,1\}}^{N\times N}$, with ’0’ on the diagonal and ’1’ in the off-diagonal positions.
@@
We propose a smoothing strategy that goes beyond simple binary categorization of matrices as positive or negative and applies a form of smoothing to the standard contrastive loss. This strategy allows nodes initially categorized as positive or negative to have values ranging from ’0’ to ’1’, indicating their degree of association with positive or negative samples, respectively.
Definition 2.2 (Smoothing process).
@@
The smoothing process $\mathcal{S}({\Pi}_{\text{pos}}^{(i,j)},\mathbf{A}^{(i)})$ generate a smooth positive matrix $\tilde{\Pi}_{\text{pos}}^{(i,j)}\in{[0,1]}^{N\times N}$ by iteratively updating binary values of $\Pi_{\text{pos}}^{(i,j)}$ based on the neighboring nodes values in the graph structure $\mathbf{A}^{(i)}$ while preserving the underlying graph structure. The corresponding smoothed negative matrix $\tilde{\Pi}_{\text{neg}}^{(i,j)}$ is then computed as $1-\tilde{\Pi}_{\text{pos}}^{(i,j)}$.
In the following, we introduce three formulations for prompting smoothing, including Taubin smoothing [15], Bilateral smoothing [16], and Diffusion-based smoothing [17]. Our first objective is to enrich both positive and negative pairs by incorporating neighborhood relationships and capturing the broader context of the nodes. Secondly, we aim to demonstrate how these enriched positive and negative sets can lead to a more effective contrastive loss.
@@
Taubin smoothing $\mathcal{S}_{T}(\mathbf{V},\mathbf{L};K,\mu,\tau)$ involves iteratively performing two stages of filtering utilized Laplacian matrix $\mathbf{L}\in\mathbb{R}^{N\times N}$ to smooth the binary matrix $\mathbf{V}\in\{0,1\}^{N\times D}$ as follows:
$\mathbf{V}^{(k+1)}=(\mathbf{I}+\tau\mathbf{L})((\mathbf{I}+\mu\mathbf{L})\mathbf{V}^{(k)})$ (2)
This process involves the combined operation of two filters, collaboratively leading to the smoothing of the input signal $\mathbf{V}$. The first filter, the negative Laplacian filter ($\mu<0$), smooths the input signal, while the second, the positive Laplacian filter $(\mathbf{I}+\tau\mathbf{L})$ ($\tau>0$), prevents oversmoothing by ensuring $\mu<-\tau$. In our approach, we employ symmetrically normalized graph Laplacian $\mathbf{L}=\mathbf{I}-\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$.
@@
Bilateral smoothing $\mathcal{S}_{B}(\mathbf{V},\mathbf{A};\sigma_{\text{spa}},\sigma_{\text{int}})$ smooths a binary matrix $\mathbf{V}$ by integrating information from nearby nodes, considering both spatial proximity and intensity similarity. Spatial proximity $d_{\text{spa}}(i,j)$ is measured using shortest path distances, while intensity similarity $d_{\text{int}}(i,j)$ is determined by evaluating the similarity in binary values between two nodes, typically quantified using metrics like the Hamming distance. The bilateral filter weight $w(i,j)$ is then computed by:
$w(i,j)=\exp\left(-\frac{d_{\text{spa}}(i,j)}{2\sigma_{\text{spa}}^{2}}-\frac{d_{\text{int}}(i,j)}{2\sigma_{\text{int}}^{2}}\right),$ (3)
where $\sigma_{\text{spa}}^{2}$ and $\sigma_{\text{int}}^{2}$ control the smoothing effects for spatial and intensity components, respectively. The smoothed value for node $v_{i}$ is computed as a weighted average of its $k$-hope neighbor nodes:
$\tilde{\textbf{v}}_{i}=\frac{\sum_{j\in\mathcal{N}_{k}(i)}w(i,j)\textbf{v}_{j}}{\sum_{j\in\mathcal{N}_{k}(i)}w(i,j)}.$ (4)
@@
Diffusion-based smoothing $\mathcal{S}_{D}(\mathbf{V},\mathbf{A};K,\eta)$ employs the diffusion equation to propagate information among nodes within a graph, effectively smoothing binary values. The process starts with the original matrix $\mathbf{V}$ as the initial condition, where each binary value serves as the initial "heat" at its respective node. The new value for each node is then iteratively updated based on the diffusion equation and the binary values of its neighbors as $\textbf{v}_{i}^{(k+1)}=\textbf{v}_{i}^{(k)}+\eta\bar{\textbf{v}}_{i}^{(k)}$, where $\bar{\textbf{v}}_{i}^{(k)}=\sum_{j\in\mathcal{N}(v_{i})}\textbf{v}_{j}^{(k)}$ is the average value of neighboring nodes, and $\eta$ is the diffusion rate applied to determine how much the binary value diffuses from one node to another.
@@
Appendix B provides a comprehensive overview of smoothing approaches, including detailed algorithms and a comparative analysis of each method. Figure 1 illustrates an example of the efficacy of the smoothing approaches. As a simple example, we take a grid graph, and randomly establish a delta function, centered on specific vertices, resulting in the creation of a binary matrix. Subsequently, we employ a variety of smoothing techniques on this binary matrix. Given the uniform neighborhood structure of the grid, the resulting output exhibits a Gaussian-like distribution, which its center aligned to the initial vertex. However, the varied values in the smoothed matrix are indicative of the distinct strategies employed in the smoothing process.
@@
Figure 2: In the general context of conventional contrastive learning approaches, for every anchor node $v_{4}^{(i)}$ in $\mathcal{G}^{(i)}$, a corresponding positive node $v_{4}^{(j)}$ exists in $\mathcal{G}^{(j)}$, with all other node pairs being negative (left image). Smoothing techniques, which leverages the geometry of graph $\mathcal{G}^{(j)}$, effectively extract neighboring node information of node $v_{4}^{(j)}$ and generate smoothed positive and negative pairs matrices $\tilde{\Pi}_{\text{pos}}^{(i,j)}$ and $\tilde{\Pi}_{\text{neg}}^{(i,j)}$ (right image).
2.5 Smoothness promoting in positive and negative sets